OpenStudy (anonymous):

2x+1 over 3=x+6 over 2 Solve for x

OpenStudy (anonymous):

step 1: cross multiply Have you heard of that before?

OpenStudy (anonymous):

with the denominator?

jigglypuff314 (jigglypuff314):

first multiply both sides by 2, then multiply both sides by 3 so \[\frac{ 2x+1 }{ 3 } = \frac{ x+6 }{ 2 } \rightarrow \frac{ 2(2x+1) }{ 3 }= x+6 \rightarrow 2(2x+1) = 3(x+6)\] then distribute so 2(2x+1) = 3(x+6) 2(2x)+2(1) = 3(x)+3(6) then simplify 4x + 2 = 3x + 18 then subtract 2 from both sides -2 -2 4x + 0 = 3x + 16 then subtract 3x from both sides -3x -3x which will give you your answer